# NuGraph2 with Explainability: Post-hoc Explanations for Geometric Neural Network Predictions

### v1

@@ 1 Introduction
As the usage for artificial intelligence (AI) becomes more pervasive across all domains of science and society, the need for explainability of AI methods, used e.g. to make a scientific discovery or to drive certain decisions in medical procedures, becomes more urgent than ever. Such methods are often regarded as black-boxes, whose power remains obscure even to the understanding of experts. The lack of explainability, thus, makes it, on the one hand, difficult to communicate and accept their outcome, and on the other hand, prevents experts from learning which features in their data played a significant role. This can leave experts’ unaware of potential unknown correlations or, worse, of unexpected shifts in the data distributions.
@@
Past work in the domain of AI explainability has brought significant advancements for convolutional neural networks (CNNs), with the development of tools and techniques that can be broadly applied to different applications (see Sec. 2.2). In the past years, Graph Neural Networks (GNNs) have emerged as a powerful generalization of CNNs that operate on data layouts that are not necessarily image-like but are rather described as a graph. This graph data structure consists of a set of nodes and a set of edges connecting them. Explainability tools for GNNs are less widely adopted than those for CNNs and, due to the larger flexibility of graph-based data structures, they often are less generalizable. In this paper, we explore explainability approaches in the context of a specific GNN for High Energy Physics (HEP) experiments, aiming at providing insights, ideas, and methods that can be valuable for the larger GNN context. There exist a general set of questions one can ask of neural network explainability methods, and we narrow these as incremental steps towards a wider understanding of GNNs. We are interested in probing the network behavior in terms of information flow and learned notions. In this paper we address a set of specific questions, and test or develop tools for extracting this information from the NuGraph GNN [1], such as: "Why did the network get a prediction wrong or correct?", "What information is captured in the network’s internal representation, and how does the training process influence it?", "How does the information flow across the graph edges, and what is the role of nodes with different edge degree?", "What is the role of network hyperparameters, such as the number of message-passing iterations?" Ultimately, we aim at addressing more fundamental questions, including: "Does it correctly encode physical laws present in the training data without being explicitly told?", "How does the network “conceptualize“ different features of inputs?", "How does the network utilize positional information versus edge information?", "How does the network, if at all, learn derived graph-wide features?". The work presented here represents a fundamental step in the direction towards such ultimate questions. We present multiple techniques that aim to address overlapping interpretations of these questions.
@@
We also make a distinction between “interpretable“ and “explainable“ results. While these terms are often interchangeable and their definitions fluid, it is generally considered that these terms do not describe exactly analogous ideas. For the sake of this paper, we adopt the definition of interpretable put forward by Doshi-Valez and Kim [2] - "ability to explain or to present in understandable terms to a human". And while a large part of interpretable and explainability works aim to make a ‘white-box‘ network as a proxy [3], we follow the narrower definition of "explainability" give by Gunning and Aha [4], where explainability is a set of methods used after the network is trained. We aim to give attributions for prediction results from the network, not to create a transparent (interpretable) network. Our methods are all used post-hoc, and examine the de-facto way the network has trained.
@@ 2.1 Case Study Application - NuGraph
NuGraph2 [1] is a GNN for the classification of energy deposits from liquid argon time projection chamber (LArTPC) neutrino experiments. LArTPC detectors record the trajectories of charged particles, produced by the interaction of neutrinos with the argon or by cosmic ray backgrounds. These particle ionize the argon along their path and, due to the uniform electric field applied to the detector, the ionization electrons drift towards the anode plane. Here they are detected in the form of electric pulses on sense wire arranged in different planes, typically three. Each plane records a two-dimensional trajectory in wire space and time, and due to the different direction of the wires, when these are combined across planes the full three-dimensional trajectory can be recovered. After the electric signals on the wires are pre-processed, pulses are fit as one or more Gaussian distribution, where each distribution represents the energy deposition of a single particle, or "hit". Hits contain information such as the wire and time of detection, the integral of the Gaussian (proportional to the energy that was deposited by the particle), and its width.
@@
NuGraph2 is trained on the MicroBooNE public data sets [6, 7], which feature simulated neutrino interactions from the Booster Neutrino Beam overlaid on top of off-beam data collected with the MicroBooNE detector [8]. Results presented in this paper are also evaluated on these data sets. A more complete description of these data sets can be found in [9]. NuGraph2 achieves an overall accuracy of 98% and 95% for the filter and semantic decoders, respectively. The performance of the semantic decoder is not uniform across the different categories, so that the decoder performs better on the most represented category int the training dataset (MIP) and worst on the least represented category, Michel. Usually, mistakes take place when the network is unsure about the classification of a group of hits, resulting in a mixed classification of that group across multiple categories. In rare circumstances, large scale failures can also happen, where most of the hits in an event (or in a particle in the event) are incorrectly classified. Understanding the origin of these mistakes is a fundamental step that can drive developments aiming at further improvements of the model.
@@ 2.2 Literature Review
The study of explainablity in graph neural networks is not a novel field. Taxonomy studies [10] [11] breaks down the majority of GNN-focused methods into two broad categories for the analysis for trained networks - explanations for single inference result and explanations for the model itself. Some notable methods include ‘GNNExplainer‘ [12] and ‘PGExplainer‘ [13] which use masking methods to determine which parts of the graph most influence a nodes’ or edges’ inference results, ‘SubgraphX‘ [14], which uses a tree search method with Shapley evaluation metrics, and ‘GraphLIME‘ [15] which adapts the popular method ‘LIME‘ [16] to GNNs. All of these are focused on inference-level explanations. The best known whole-model explanation method is the generative ‘XGNN‘ [17] method, which endeavors to encode the "rules" of the network to produce a small class of subgraphs representing the model’s "understanding" of the input data.
@@ 3.1 Network Explanations
Much of the existing literature for post-training graph network explainability focuses on geometric situations with physical edges, or edges with some meaning relative to the nodes. By contrast, NuGraph2’s edges are all constructed algorithmically using Delaney triangulation. The process of reconstructing the tracks to align with physical constraints would require knowledge of the identity of particles leaving the tracks, and render the identification network redundant. This caused us to explore applications of existing methods that can work around the nonphysical edges, such as GNNExplainer [12].
@@
A total of 6 masks (one for each planar graph, and one for each set of connections between the planar graph the nexus layer) are initialized and jointly optimized over graphs $G$ through the entropy minimization equation $\mathcal{L}=H(Y)-H(Y|G=G_{S})$. In our case, $G_{S}$ is defined as the subgraph with all 6 masks applied. Afterwards, we then examine the individual subgraphs constructed using the ground truth of our data. Through this we aim to examine two classes of questions.
@@
The above approaches have two main limitations. The first is that the gradient based optimization method central to the discovery of important subgraphs in GNNExplainer cannot reach satisfactory minima in NuGraph2’s graphs. This is shown in table 1, where none of the methods reach a sufficient Fidelity± score to declare the subgraphs representative of the graph. In fact, having both Fidelity+ and Fidelity- values close to one indicates that all graph features are considered important for the observed result. Since the method is not able to identify which aspects are less relevant, explanations provided lack in conciseness and are ultimately not providing useful insight related to the specific outcome (Fig. 2). We hypothesize that methods that define masks describing relevant sub-graphs for the observed outcome struggle with NuGraph2 due to the a large degree of redundancy in the edge connections. We verify this hypothesis in Section 3.2.
The second limitation has to do with the combination of the base method and the information we want to extract. GNNExplainer strives to make subgraphs that explain small groups of nodes or edges in the graph, and while it can be applied to whole graphs, it cannot capture a comprehensive explanation of the whole network’s "understanding" or encoding of the problem as a whole. GNNExplainer, in the case of NuGraph2, is limited to semi-local explanations of singular graph, and as a consequence relies on the practitioner to find examples of certain kind of behaviors seen in the data, with no guarantee the inferences made from a singular graph extend to the problem as a whole. To address this limitation, in Sections 3.3 and 3.4 we investigate how the network is able to distinguish between different general concepts.
@@ 3.2 Graph Rewiring
To further address the question of "How many edges are enough?" and investigate the inductive biases added by the Delaunay triangulation algorithm, we conduct graph rewiring and edge pruning. Delauney triangulation is used to create a graph from the point cloud data. This wiring algorithm subdivides the convex hull of a set of points on a plane into triangles whose circumcircles (i.e. circles that pass through all the three vertices forming the triangle) don’t contain any of the points. While connecting point clouds using Delauney triangulation does not immediately lead to physically interpretable connections, it nevertheless carries learning biases due to the algorithm definition, such as:
  • Locality. Nearest-neighbors are by construction a preferred connection path with Delauney triangulation. Long-range dependencies are present but less frequently created. In NuGraph2, isolated nodes act as hubs of long-range connections. This implies that information across distant nodes needs to travel through several message passing iterations or through hub connections.
@@ 3.3 Latent Space
Large change in the quality of clusters between steps allow to see changes in the latent representation of labels and how well the latent space maps to labels - indicating when the encoding is stable. The order of cluster formation for labels can similarly show us the successes and short-comings of the classification method. Clearly, the network identifies MIP and background nodes first, as they are the most represented category in the dataset. On the other hand, it is not until the last two iterations that categories with smaller population such as shower and diffuse stand out as a cluster. The fact that there are diminishing improvements from the fourth to the last iteration supports the choice of five total iterations.
@@
Notably, because of the 2D representation of the embeddings in figure 7, there is not perfect agreement between the visualization of the cluster and the silhouette scores. The consistent downgrade in cluster clarity between the encoder stop and the first message passing step indicates that the immediate neighbors of a given node are not useful for classification, and the non-local nature of message passing is important to quality predictions. The large jump in cluster separation between 3 and 5 also supports the choice of 5 iterations as nominal configuration in NuGraph2 [1]. Despite the relatively high number of message passing steps used, this is the optimal number of messages for this network. It does also show the failure of the network to encode Michel electrons, shown by the decrease in cluster clarity. Logically, this implies that the fewer number of Michel electrons are not weighted sufficiently within the dataset to overcome the amount other classes updating the node’s hidden representation.
@@ 3.4 Decoder Probes
These concepts represent emergent properties of the data that can be learned by the network. As an analogy, this can be thought of testing the ability of a convolutional network trained on images of zebras to recognize stripes as a generalized concept. This work is inspired by the notion of "Testing with Concept Activation Vectors" [23]. The key difference is that we do not use examples to construct a subset which represents our concept, rather, we produce a custom loss to measure how accurately selected concepts are learned. Using this method of decoder probes, this idea can be extended across other concepts that can be represented with a loss function. This means that this approach can be used to evaluate to which extent the network learns arbitrary concepts (which can be directly related to the primary learning task of the training, but not necessarily), thus effectively providing a tool to evaluate whether the network is gaining knowledge about underlying domain concepts. Such tool could be used e.g. to evaluate whether a given network shows properties similar to a foundation model.
@@
Three test concepts were selected; two concepts represent relationships between specific types of particle hits within the graph, and one concept aims at showing how one semantic class is learned. All utilize a modification of cross entropy loss focusing on a certain semantic class behavior.
@@
    Tracks are the "easiest" concept to identify among the concepts tested. This is judged by seeing the number of network steps needed before the probe losses plateau. We can see this in figure 9, where losses are roughly stabilized by network step two. This implies there is very little left to encode in the latent space regarding the relationship between HIP and MIP particles in the data the network has access to, and additional layers or operations applied to the data would not provide additional accuracy. This can lead to the assumption that as this is the easiest concept to identify, it could also be the easiest concept overlearn as well.
@@
    The use of non-local information from the message passing steps helps both "difficult" and "easy" concepts, but not to the same degree. There is no place in the network where we see a decrease in concept accuracy from the probes, nor a spot where the probes plateau only to grok the concept afterwards. This implies there is not a number of hops where the nodes in the graph become actively dissimilar or unrelated for any of the tested concepts, decreasing the accuracy of the concept probes, nor is there a minimum number of hops needed to improve the concept accuracy. While this is most likely not applicable to all problems, it is important to recognize that the construction of our problem results in a situation where non-local information, at least to 5 hops away, does not decrease the accuracy of these concepts.
@@ 5 Acknowledgments
The authors would like to thank V Hewes for their assistance in coding NuGraph extensions and understanding the intricacies of the original network’s implementation, and Burt Holtzman and Maria Acosta for their management of computing resources used for running a large amount of the analysis included in this work. We acknowledge the MicroBooNE Collaboration for making publicly available the data sets [7, 6] employed in this work.

### v2

@@ 1 Introduction
As the usage for artificial intelligence (AI) becomes more pervasive across all domains of science and society, the need for explainability of AI methods, used e.g. to make a scientific discovery or to drive certain decisions in medical procedures, becomes more urgent than ever. Such methods are often regarded as black-boxes, whose power remains obscure even to the understanding of experts. The lack of explainability, thus, makes it, on the one hand, difficult to communicate and accept their outcome, and on the other hand, prevents experts from learning which features in their data played a significant role. This can leave experts unaware of potential unknown correlations or, worse, of unexpected shifts in the data distributions.
@@
Past work in the domain of AI explainability has brought significant advancements for convolutional neural networks (CNNs), with the development of tools and techniques that can be broadly applied to different applications (see Sec. 2.2). In the past years, Graph Neural Networks (GNNs) have emerged as a powerful generalization of CNNs that operate on data layouts that are not necessarily image-like but are rather described as a graph. This graph data structure consists of a set of nodes and a set of edges connecting them. Explainability tools for GNNs are less widely adopted than those for CNNs and, due to the larger flexibility of graph-based data structures, they often are less generalizable. In this paper, we explore explainability approaches in the context of a specific GNN for High Energy Physics (HEP) experiments, aiming at providing insights, ideas, and methods that can be valuable for the larger GNN context. There exist a general set of questions one can ask of neural network explainability methods, and we narrow these as incremental steps towards a wider understanding of GNNs. We are interested in probing the network behavior in terms of information flow and learned notions. In this paper we address a set of specific questions, and test or develop tools for extracting this information from the NuGraph GNN [1], such as: "Why did the network get a prediction wrong or correct?", "What information is captured in the network’s internal representation, and how does the training process influence it?", "How does the information flow across the graph edges, and what is the role of nodes with different edge degree?", "What is the role of network hyperparameters, such as the number of message-passing iterations?" Ultimately, we aim to address more fundamental questions, including: "Does it correctly encode physical laws present in the training data without being explicitly told?", "How does the network “conceptualize“ different features of inputs?", "How does the network utilize positional information versus edge information?", "How does the network, if at all, learn derived graph-wide features?". The work presented here represents a fundamental step in the direction towards such ultimate questions. We present multiple techniques that aim to address overlapping interpretations of these questions.
@@
We also make a distinction between “interpretable“ and “explainable“ results. While these terms are often interchangeable and their definitions fluid, it is generally considered that these terms do not describe exactly analogous ideas. For the sake of this paper, we adopt the definition of interpretable put forward by Doshi-Valez and Kim [2] - "ability to explain or to present in understandable terms to a human". While a large part of interpretable and explainability works aim to make a ‘white-box‘ network as a proxy [3], we follow the narrower definition of "explainability" give by Gunning and Aha [4], where explainability is a set of methods used after the network is trained. We aim to give attributions for prediction results from the network, not to create a transparent (interpretable) network. Our methods are all used post-hoc, and examine the de-facto way the network has trained.
@@ 2.1 Case Study Application - NuGraph
NuGraph2 [1] is a GNN for the classification of energy deposits from liquid argon time projection chamber (LArTPC) neutrino experiments. LArTPC detectors record the trajectories of charged particles, produced by the interaction of neutrinos with the argon or by cosmic ray backgrounds. These particle ionize the argon along their path and, due to the uniform electric field applied to the detector, the ionization electrons drift towards the anode plane. Here they are detected in the form of electric pulses on sense wires arranged in different planes, typically three. Each plane records a two-dimensional trajectory in wire space and time, and due to the different direction of the wires, when these are combined across planes the full three-dimensional trajectory can be recovered. After the electric signals on the wires are pre-processed, pulses are fit as one or more Gaussian distribution, where each distribution represents the energy deposition of a single particle, or "hit". Hits contain information such as the wire and time of detection, the integral of the Gaussian (proportional to the energy that was deposited by the particle), and its width.
@@
NuGraph2 is trained on the MicroBooNE public data sets [6, 7], which feature simulated neutrino interactions from the Booster Neutrino Beam overlaid on top of off-beam data collected with the MicroBooNE detector [8]. Results presented in this paper are also evaluated on these data sets. A more complete description of these data sets can be found in [9]. NuGraph2 achieves an overall accuracy of 98% and 95% for the filter and semantic decoders, respectively. The performance of the semantic decoder is not uniform across the different categories, so that the decoder performs better on the most represented category in the training dataset (MIP) and worst on the least represented category, Michel. Usually, mistakes take place when the network is unsure about the classification of a group of hits, resulting in a mixed classification of that group across multiple categories. In rare circumstances, large scale failures can also happen, where most of the hits in an event (or in a particle in the event) are incorrectly classified. Understanding the origin of these mistakes is a fundamental step that can drive developments aiming at further improvements of the model.
@@ 2.2 Literature Review
The study of explainablity in graph neural networks is not a novel field. Taxonomy studies [10] [11] breaks down the majority of GNN-focused methods into two broad categories for the analysis for trained networks - explanations for single inference result and explanations for the model itself. Some notable methods include ‘GNNExplainer‘ [12] and ‘PGExplainer‘ [13] which use masking methods to determine which parts of the graph most influence a node’s inference results, ‘SubgraphX‘ [14], which uses a tree search method with Shapley evaluation metrics, and ‘GraphLIME‘ [15] which adapts the popular method ‘LIME‘ [16] to GNNs. All of these are focused on inference-level explanations. The best known whole-model explanation method is the generative ‘XGNN‘ [17] method, which endeavors to encode the "rules" of the network to produce a small class of subgraphs representing the model’s "understanding" of the input data.
@@ 3.1 Network Explanations
Much of the existing literature for post-training graph network explainability focuses on geometric situations with physical edges, or edges with some meaning relative to the nodes. By contrast, NuGraph2’s edges are all constructed algorithmically using Delaunay triangulation. The process of reconstructing the tracks to align with physical constraints would require knowledge of the identity of particles leaving the tracks, and render the identification network redundant. This caused us to explore applications of existing methods that can work around the nonphysical edges, such as GNNExplainer [12].
@@
A total of 6 masks (one for each planar graph, and one for each set of connections between the planar graph and the nexus layer) are initialized and jointly optimized over graphs $G$ through the entropy minimization equation $\mathcal{L}=H(Y)-H(Y|G=G_{S})$. In our case, $G_{S}$ is defined as the subgraph with all 6 masks applied. Afterwards, we then examine the individual subgraphs constructed using the ground truth of our data. Through this we aim to examine two classes of questions.
@@
The above approaches have two main limitations. The first is that the gradient based optimization method central to the discovery of important subgraphs in GNNExplainer cannot reach satisfactory minima in NuGraph2’s graphs. This is shown in table 1, where none of the methods reach a sufficient Fidelity± score to declare the subgraphs representative of the graph. In fact, having both Fidelity+ and Fidelity- values close to one indicates that all graph features are considered important for the observed result. Since the method is not able to identify which aspects are less relevant, explanations provided lack in conciseness and are ultimately not providing useful insight related to the specific outcome (Fig. 2). We hypothesize that methods that define masks describing relevant sub-graphs for the observed outcome struggle with NuGraph2 due to the large degree of redundancy in the edge connections. We verify this hypothesis in Section 3.2.
The second limitation has to do with the combination of the base method and the information we want to extract. GNNExplainer strives to make subgraphs that explain small groups of nodes or edges in the graph, and while it can be applied to whole graphs, it cannot capture a comprehensive explanation of the whole network’s "understanding" or encoding of the problem as a whole. GNNExplainer, in the case of NuGraph2, is limited to semi-local explanations of singular graphs, and as a consequence relies on the practitioner to find examples of certain kind of behaviors seen in the data, with no guarantee the inferences made from a singular graph extend to the problem as a whole. To address this limitation, in Sections 3.3 and 3.4 we investigate how the network is able to distinguish between different general concepts.
@@ 3.2 Graph Rewiring
To further address the question of "How many edges are enough?" and investigate the inductive biases added by the Delaunay triangulation algorithm, we conduct graph rewiring and edge pruning. Delaunay triangulation is used to create a graph from the point cloud data. This wiring algorithm subdivides the convex hull of a set of points on a plane into triangles whose circumcircles (i.e. circles that pass through all the three vertices forming the triangle) don’t contain any of the points. While connecting point clouds using Delaunay triangulation does not immediately lead to physically interpretable connections, it nevertheless carries learning biases due to the algorithm definition, such as:
  • Locality. Nearest-neighbors are by construction a preferred connection path with Delaunay triangulation. Long-range dependencies are present but less frequently created. In NuGraph2, isolated nodes act as hubs of long-range connections. This implies that information across distant nodes needs to travel through several message passing iterations or through hub connections.
@@ 3.3 Latent Space
Large change in the quality of clusters between steps allow us to see changes in the latent representation of labels and how well the latent space maps to labels - indicating when the encoding is stable. The order of cluster formation for labels can similarly show us the successes and short-comings of the classification method. Clearly, the network identifies MIP and background nodes first, as they are the most represented categories in the dataset. On the other hand, it is not until the last two iterations that categories with smaller populations such as shower and diffuse stand out as a cluster. The fact that there are diminishing improvements from the fourth to the last iteration supports the choice of five total iterations.
@@
Notably, because of the 2D representation of the embeddings in figure 7, there is not perfect agreement between the visualization of the cluster and the silhouette scores. The consistent downgrade in cluster clarity between the encoder stop and the first message passing step indicates that the immediate neighbors of a given node are not useful for classification, and the non-local nature of message passing is important to quality predictions. The large jump in cluster separation between 3 and 5 also supports the choice of 5 iterations as nominal configuration in NuGraph2 [1]. Despite the relatively high number of message passing steps used, this is the optimal number of messages for this network. It does also show the failure of the network to encode Michel electrons, shown by the decrease in cluster clarity. Logically, this implies that the fewer number of Michel electrons are not weighted sufficiently within the dataset to overcome the contributions from other classes updating the node’s hidden representation.
@@ 3.4 Decoder Probes
These concepts represent emergent properties of the data that can be learned by the network. As an analogy, this can be thought of testing the ability of a convolutional network trained on images of zebras to recognize stripes as a generalized concept. This work is inspired by the notion of "Testing with Concept Activation Vectors" [23]. The key difference is that we do not use examples to construct a subset which represents our concept - rather, we produce a custom loss to measure how accurately selected concepts are learned. Using this method of decoder probes, this idea can be extended across other concepts that can be represented with a loss function. This means that this approach can be used to evaluate to what extent the network learns arbitrary concepts (which can be directly related to the primary learning task of the training, but not necessarily), thus effectively providing a tool to evaluate whether the network is gaining knowledge about underlying domain concepts. Such tool could be used e.g. to evaluate whether a given network shows properties similar to a foundation model.
@@
Three test concepts were selected: two concepts represent relationships between specific types of particle hits within the graph, and one concept aims at showing how one semantic class is learned. All utilize a modification of cross entropy loss focusing on a certain semantic class behavior.
@@
    Tracks are the "easiest" concept to identify among the concepts tested. This is judged by seeing the number of network steps needed before the probe losses plateau. We can see this in figure 9, where losses are roughly stabilized by network step two. This implies there is very little left to encode in the latent space regarding the relationship between HIP and MIP particles in the data the network has access to, and additional layers or operations applied to the data would not provide additional accuracy. This can lead to the assumption that as this is the easiest concept to identify, it could also be the easiest concept to overlearn as well.
@@
    The use of non-local information from the message passing steps helps both "difficult" and "easy" concepts, but not to the same degree. There is no place in the network where we see a decrease in concept accuracy from the probes, nor a spot where the probes plateau only to grasp the concept afterwards. This implies there is not a number of hops where the nodes in the graph become actively dissimilar or unrelated for any of the tested concepts, decreasing the accuracy of the concept probes, nor is there a minimum number of hops needed to improve the concept accuracy. While this is most likely not applicable to all problems, it is important to recognize that the construction of our problem results in a situation where non-local information, at least to 5 hops away, does not decrease the accuracy of these concepts.
@@ 5 Acknowledgments
The authors would like to thank Burt Holtzman and Maria Acosta for their management of computing resources used for running a large amount of the analysis included in this work. We acknowledge the MicroBooNE Collaboration for making publicly available the data sets [7, 6] employed in this work.